The total age of a woman and her son is 51 years. Three years ago, the woman was eight times as old as her son. How old is her son now?
Answer by richwmiller(17219)

w+s=51,
w=51-s
w-3=8(s-3)
w-3=8s-24
w=8s-21
8s-21=51-s
9s=72
s=8
w=43